OpenStudy (anonymous):

Please help! Described how the graph of y = (2x+2)^2 can be obtained from the graph of y = x^2.

OpenStudy (anonymous):

But how can graph 1 be obtained from graph 2? I'm not sure how this helps me. What do I do with this?

OpenStudy (anonymous):

Make different plots until u see the pattern, that way you will learn it better. For example http://www.wolframalpha.com/input/?i=Plot+%28x-1%29%5E2%2C+%28x%2B1%29%5E2%2C+x%5E2+for+x+in+-5..5

OpenStudy (anonymous):

The pattern I see is that when x is squared, it is a parabola. Is there another pattern I should be looking for?

jimthompson5910 (jim_thompson5910):

y = (2x+2)^2 y = (2(x+1))^2 y = 2^2*(x+1)^2 y = 4(x+1)^2 Compare this with y = a(x-h)^2 + k

OpenStudy (anonymous):

All of those equations are the same parabola, right?

jimthompson5910 (jim_thompson5910):

they are

OpenStudy (anonymous):

I'm not sure what I'm supposed to do with that.

jimthompson5910 (jim_thompson5910):

Say for example you go from y = x^2 to y = (x-1)^2 You're shifting it 1 unit to the right

jimthompson5910 (jim_thompson5910):

if you can get it into the form y = a(x-h)^2 + k, you can determine how to shift and how to stretch/compress it

OpenStudy (anonymous):

But adding k would be determining where on the y-axis the vertex would be. So why would I want to change the y-axis vertex?

jimthompson5910 (jim_thompson5910):

you can easily make k = 0 to make sure it doesn't shift up/down

OpenStudy (anonymous):

Oh! I think I get it now!

OpenStudy (anonymous):

Kinda... So basically the answer to this question would be explaining how to evolve y = (2x+2)^2 into the y = a(x-h)^2 + k format. Am I anywhere close to what you've been trying to explain?

jimthompson5910 (jim_thompson5910):

yes you are you turn y = (2x+2)^2 into y = 4(x+1)^2 which is in y = a(x-h)^2 + k form where a = 4, h = -1 and k = 0

jimthompson5910 (jim_thompson5910):

since h = -1 , you shift y = x^2 1 unit to the left you don't shift up or down because k = 0

jimthompson5910 (jim_thompson5910):

a = 4 means that you stretch everything vertically by a factor of 4

OpenStudy (anonymous):

Thank you. I think I have a better understanding of this material. Now I just have to put it into words on the page. :P

jimthompson5910 (jim_thompson5910):

yw